# Towards Better Open-Ended Text Generation: A Multicriteria Evaluation Framework

### v2

@@ Abstract
Open-ended text generation has become a prominent task in natural language processing due to the rise of powerful (large) language models. However, evaluating the quality of these models and the employed decoding strategies remains challenging because of trade-offs among widely used metrics such as coherence, diversity, and perplexity. Decoding methods often excel in some metrics while underperforming in others, complicating the establishment of a clear ranking. In this paper, we present novel ranking strategies within this multicriteria framework. Specifically, we employ benchmarking approaches based on partial orderings and present a new summary metric designed to balance existing automatic indicators, providing a more holistic evaluation of text generation quality. Our experiments demonstrate that the proposed methods offer a robust way to compare decoding strategies, and serve as valuable tools in guiding model selection for open-ended text generation tasks. Finally, we suggest future directions for improving evaluation methodologies in text generation. Our codebase, datasets, and models are publicly available. [^1] https://github.com/YecanLee/2BeOETG
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in generating coherent and contextually appropriate text across various domains. However, evaluating the quality of their generated outputs remains a substantial challenge. Traditional evaluation methods often rely on human judgment or automated metrics, each capturing different aspects of text quality but none providing a comprehensive assessment.
@@
Figure 1: We present a multicriteria evaluation framework for benchmarking models and decoding strategies, i.e., decoding methods. We distinguish two scenarios for benchmarking (Section 1) and two ranking objectives (Section 4), giving rise to three use-case tailored, distinct methods (sections 4.14.2 and 5).
Decoding methods frequently exhibit trade-offs, excelling in some metrics while underperforming in others, complicating the establishment of a consistent ranking among them. Metrics such as MAUVE (Pillutla et al., 2021), coherence (Su et al., 2022), diversity, and generation perplexity (Jelinek et al., 2005) are commonly used to assess the quality of generated text and the generation confidence. While MAUVE measures the alignment between generated and human-like text distributions, coherence evaluates logical flow, diversity ensures varied outputs, and perplexity assesses the predictability of text. While each metric provides valuable insights, considering them in isolation offers an incomplete perspectives on the overall generation quality. Human evaluation, though often considered the gold standard, is resource-intensive and subjective, making it impractical for large-scale assessments of generated texts. In this work, we focus on automatic evaluation of generated text. The challenge lies in balancing multiple automatic (and often contradictory) metrics to create a comprehensive evaluation framework that can reliably compare different models and decoding strategies – collectively referred to as decoding methods in this paper. More LLMs and decoding strategies imply that the number of such decoding methods grows strongly. This development renders reliable benchmarking methods for comparing them even more important.
In practice, two prototypical examples of benchmarking scenarios can be distinguished, see Fig. 2.
Scenario 1: Ranking Firstly, consider a practitioner using open-ended text generation for a very specific task, e.g., a customer support chatbot. This practitioner might primarily be interested in simply knowing which decoding method works best for this particular scenario and may additionally be keen to know which method comes second, third, etc. This motivation renders metric information about the methods’ performances a means to an end. Thus, a ordinal ranking of methods will do, see Figure 2 (top).
Scenario 2: Metric Assessment Secondly, think of a researcher who is interested in designing new decoding methods (i.e., either novel models or novel decoding strategies or both). For such a researcher, it is of utmost importance to know how much better one method is compared to another one, i.e., have metric information, see Figure 2 (bottom). Knowing about how existing decoding methods perform on different tasks might help them to derive new methods.
In reality, there are of course many more scenarios in between these two prototypical cases. In this work, we propose multiple benchmarking methods along this spectrum. What unites them is their ability to facilitate multiple criteria. This way, we address the following issues.
@@
Research Gap: When evaluating decoding methods based on multiple criteria (such as metrics or tasks) across several test problems (like different datasets), a method may excel in one area while lagging in another. It is still understood as an open problem of how to aggreate such multicriteria evaluation results in different scenarios, see Figure 1 and 2. Existing approaches comprise the Pareto front – all decoding methods that are undominated across all criteria – or weighted sums of the criteria. While the former wastes a lot of information, see also Section 4, the latter relies arbitrarily on the selected weights. This paper offers alternatives. In particular, we address the following questions.
Research Questions
  1. 1.
    RQ1: Can we exploit novel statistical methodologies for partial orders to establish multicriteria rankings that potentially allow for incomparability? (Scenario 1: Ranking)
  2. 2.
    RQ2: Can we aggregate multiple automatic evaluation metrics in a meaningful and statistically valid way? (Scenario 2: Metric Assessment)
We address RQ1 in Section 4 and RQ2 in Section 5 by proposing appropriate aggregation methods, see also Figure 1. What is more, we apply all these methods on 1.8 million stories generated by six state-of-the-art models on real world datasets, see Section 3 (setup) and Section 6 (results).
Figure 2: Illustration of ordinal ranking (top) and metric comparison (below) of decoding methods (Scenarios 1 and 2, respectively). Top: Distances are not meaningful, lower three methods are incomparable. Bottom: Distances are meaningful, all methods are comparable.
@@ 2 Related Work
Benchmarks are the backbone of any kind of applied machine learning research Zhang and Hardt (2024a); Shirali et al. (2023); Ott et al. (2022); Zhang et al. (2020); Thiyagalingam et al. (2022); Roelofs et al. (2019); Vanschoren et al. (2014). On the one hand, they help practitioners make informed choices among different methods for a given problem. On the other hand, newly proposed methods typically need to demonstrate their benefits through such benchmark studies, see for instance Meyer et al. (2003); Hothorn et al. (2005); Eugster et al. (2012); Mersmann et al. (2015). In the last two decades or so, detecting statistically significant differences between methods has attracted increasing attention, recognizing the fact that problems (like datasets) in a benchmark suite are only samples of a population that needs to be defined, see Demšar (2006); García and Herrera (2008); García et al. (2010); Benavoli et al. (2017) for some pioneering works.
@@
In recent years, the focus has shifted towards multi-criteria and multi-task benchmarking problems Cruz et al. (2024); Zhang and Hardt (2024b); Kohli et al. (2024); Jansen et al. (2024, 2023b, 2023a); Rodemann and Blocher (2024); Blocher et al. (2024). In a multitude of domains, there are several criteria with respect to which methods need to be compared. Classical examples include runtime and accuracy in predictive machine learning Koch et al. (2015); Jansen et al. (2024) or performance (minimal function value achieved in a fixed time budget) and speed (time required to achieve a given function value) in optimization Schneider et al. (2018), to name only a few.
@@
This shift towards multi-citeria/task benchmarking is particularly prevalent in natural language processing (NLP). Modern generalist models, like LLMs, are evaluated with respect to a multitude of metrics, owing to their breadth and depth of capabilities, see e.g., Wei et al. (2024); Liu et al. (2025). As we can only improve upon what we measure, multi-criteria benchmarking has moved to the forefront of scientific inquiry. By assessing models across diverse tasks – ranging from reasoning and comprehension to creativity and ethical considerations – researchers gain a more holistic understanding of their strengths and limitations Chiang et al. (2024). Such comprehensive evaluation frameworks not only drive advancements in model performance but also ensure alignment with real-world applications and ethical standards Liu et al. (2023); Ji et al. (2023); Terry et al. (2023); Rodemann et al. (2025). As a result, multi-criteria benchmarking has become an essential pillar in the evolution of AI, guiding both theoretical progress and practical deployment.
Decoding methods for open-ended text generation are no exception. Several metrics to evaluate the quality of decoding strategies have been proposed and discussed in recent years Alihosseini et al. (2019); Pillutla et al. (2021); Celikyilmaz et al. (2021); Su and Xu (2022); Su et al. (2022); Gao et al. (2022); Becker et al. (2024); Garces-Arias et al. (2024a, 2025). Diversity, MAUVE, coherence, and generation perplexity are among the most popular metrics. Diversity measures lexical variation using $n$-gram repetition rates, with higher scores indicating less repetition. MAUVE is a distribution similarity metric between generated and reference text. Coherence is defined as the averaged log-likelihood of the generated text conditioned on the prompt. Finally, generation perplexity Jelinek et al. (2005) measures the predictability of the generated text under the language model; lower perplexity indicates that the text is more likely according to the model’s own probability distribution.
This multitude of quality metrics naturally gives rise to the question of how to aggregate them, i.e., how to compare decoding methods holistically, taking into account multiple dimensions of text quality. It is self-evident that focusing on only one metric has obvious shortcomings. For instance, exclusively optimizing coherence will favor decoding methods with only moderate to low penalties for already selected tokens. This can lead to repetitive and uncreative text generation; a problem often referred to as degeneration Lee et al. (2022); Holtzman et al. (2019). On the other hand, focusing solely on diversity will eventually result in incoherent text only slightly – if at all – related to the prompt.
In this work, we address exactly the problem of how to benchmark decoding strategies with respect to multiple and potentially contradicting criteria. In particular, we offer a fresh perspective on this problem, adopting recent developments in the theory of depth functions and order theory, see Section 4. Moreover, we propose a novel summary metric to balance multiple metrics, see Section 5.
3 Experimental Setup
@@
Deepseek Wikitext Coherence Beam search Beam width {3, 5, 10, 15, 20, 50} 6 × 5261 × 6 = 189,396
@@
Llama3 Sampling with temperature Temperature {0.1, 0.3, 0.5, 0.7, 0.9, 1.0} 6 × 5261 × 6 = 189,396
Mistralv03 Top $k$ sampling $k$1, 3, 5, 10, 15, 20, 50} 6 x 5261 x 7 = 220,962
Qwen2 Top $p$ (nucleus) sampling $p$0.6, 0.7, 0.8, 0.9, 0.95} 6 × 5261 × 5 = 157,830
@@
Table 1: Experimental setup. The number of computations is calculated based on the number of models, prompts, and hyperparameter values.
@@
We evaluate six model architectures that generated 1.8 Mio stories from prompts sourced from three distinct datasets, utilizing five decoding strategies across 59 hyperparameter configurations. To measure the quality of the generated text, we employed three widely used automatic metrics.
3.1 Models
@@
We employ GPT2-XL (1.5B parameters) by Radford et al. (2019), Mistral 7B v0.3 (Jiang et al., 2023, 2024), Llama 3.1 8B (Dubey et al., 2024), Deepseek 7B (DeepSeek-AI et al., 2024), Qwen 2 7B (Yang et al., 2024), and Falcon 2 11B (Malartic et al., 2024).
3.2 Evaluation Metrics
Building upon Su and Collier (2023), we employ various automatic metrics to assess the quality of the generated texts: Diversity, Coherence, and Generation Perplexity (Jelinek et al., 2005; Holtzman et al., 2019). Additional metrics like the MAUVE score measure the distributional similarity between generated texts and reference texts. However, since MAUVE relies solely on aggregated data and requires reference texts, we exclude it from our evaluations in this study. A detailed overview of their definitions is provided in Section A.1.
3.3 Datasets
@@
We evaluate our methods across three domains for open-ended text generation: news, Wikipedia articles, and stories. Specifically, we use 2,000 articles from Wikinews for the news domain; 1,314 articles from the WikiText-103 dataset (Merity et al., 2016) for the Wikipedia domain; and 1,947 examples from the Project Gutenberg split of the BookCorpus (Zhu et al., 2015) for the story domain. Each example consists of a prompt and a gold reference (i.e., a human-generated continuation) for evaluation purposes.
Furthermore, we utilize the dataset provided by Garces-Arias et al. (2024a), which includes over 2.2 million generated continuations (with a maximum length of 256 tokens) for each prompt, along with aggregated metrics such as coherence, diversity, and MAUVE. We extend this dataset by computing sentence-level evaluations and incorporating generation perplexity, as described above in Section 3.2.
3.4 Decoding Strategies and Hyperparameters
@@
For contrastive search (Su et al., 2022), we evaluate combinations of $\alpha\in\{0.2,0.4,0.6,0.8,1.0\}$ and $k\in\{1,3,5,10,15,20,50\}$, while for beam search (Freitag and Al-Onaizan, 2017), we consider beam width $\in\{3,5,10,15,20,50\}$. For sampling with temperature (Ackley et al., 1985), we consider temperature $\in\{0.1,0.3,0.5,0.7,0.9,1.0\}$, for top-$k$ sampling (Fan et al., 2018), we use $k\in\{1,3,5,10,15,20,50\}$ and for top-$p$ (nucleus) sampling (Holtzman et al., 2019) we evaluate $p\in\{0.6,0.7,0.8,0.9,0.95\}$, for a total of 59 hyperparameter combinations, totaling to 1.8 Mio generated stories and thus 1.8 Mio evaluations that we employ for benchmarking. Further details are presented in Table 1.
4 Benchmarking utilizing Order Structure

### v3

@@ Abstract
Open-ended text generation has become a prominent task in natural language processing due to the rise of powerful (large) language models. However, evaluating the quality of these models and the employed decoding strategies remains challenging due to trade-offs among widely used metrics such as coherence, diversity, and perplexity. This paper addresses the specific problem of multicriteria evaluation for open-ended text generation, proposing novel methods for both relative and absolute rankings of decoding methods. Specifically, we employ benchmarking approaches based on partial orderings and present a new summary metric to balance existing automatic indicators, providing a more holistic evaluation of text generation quality. Our experiments demonstrate that the proposed approaches offer a robust way to compare decoding strategies and serve as valuable tools to guide model selection for open-ended text generation tasks. We suggest future directions for improving evaluation methodologies in text generation and make our code, datasets, and models publicly available. [^1] https://github.com/YecanLee/2BeOETG
@@ 1 Introduction
Large language models (LLMs, e.g., Dubey et al., 2024; Yang et al., 2024) have demonstrated remarkable capabilities in generating coherent and contextually appropriate text across diverse domains. However, the quality of LLM outputs is fundamentally determined not only by the underlying model architecture but also by the decoding strategies employed during inference—the algorithms that transform the model’s output probability distributions into actual text sequences. As the landscape of both LLMs and decoding strategies continues to expand rapidly, the need for robust evaluation frameworks has become increasingly critical (Wiher et al., 2022; Garces-Arias et al., 2025).
@@
Figure 1: Multicriteria evaluation framework for benchmarking models and decoding strategies, i.e., decoding methods. We distinguish two scenarios for benchmarking (§1) and two ranking objectives (§4), giving rise to three use-case tailored, distinct methods (§4.14.3 and 5).
Scope and Problem Definition.
This paper specifically addresses the challenge of multicriteria evaluation in open-ended text generation, where we must simultaneously consider multiple, often conflicting quality dimensions (Holtzman et al., 2019; Su and Xu, 2022). We focus on developing principled methods for both relative and absolute rankings of decoding methods. Our approach centers on a subset of automatic evaluation metrics—coherence, diversity, and generation perplexity—that capture fundamental trade-offs in text generation quality. While numerous other metrics exist (e.g., relevance, informativeness, style consistency), we deliberately limit our scope to these three core dimensions to establish a foundational framework that can be systematically extended.
Current evaluation approaches face remarkable limitations when assessing the quality of text generations within this multicriteria context. Traditional methods typically rely on either human judgments—considered the gold standard, but resource-intensive, and dependent on carefully designed protocols (Howcroft et al., 2020; van der Lee et al., 2021; Karpinska et al., 2021; Ruan et al., 2024)—or individual automatic metrics. While automatic metrics such as MAUVE (Pillutla et al., 2021), coherence (Su et al., 2022), diversity, and generation perplexity (Jelinek et al., 2005) provide valuable insights into specific aspects of generation quality, an isolated consideration of these measures offers only an incomplete perspective on overall performance and fails to address the fundamental multicriteria nature of the evaluation problem.
In the context of open-ended text generation, this evaluation challenge is particularly acute because decoding strategies inherently involve trade-offs between competing objectives such as coherence and diversity. A method that excels in coherence may underperform in diversity, and vice versa, making it difficult to establish consistent relative rankings among different approaches or provide meaningful absolute assessments of their quality.
The fundamental challenge addressed in this work lies in developing principled approaches for both relative and absolute multicriteria evaluation that can balance our selected subset of automatic metrics within a comprehensive framework. This enables reliable comparison of different models and decoding strategies—collectively referred to as decoding methods throughout this work (Fig. 1)—while acknowledging the inherent trade-offs between the chosen evaluation criteria. Addressing this challenge is essential for advancing the field of open-ended text generation evaluation and providing practitioners with evidence-based guidance for selecting optimal decoding methods within the multicriteria landscape we define.
Research Gap.
@@
When evaluating decoding methods based on multiple quality criteria in several scenarios (i.e., datasets), a method may excel in one area while lagging in another. Aggregating such multicriteria evaluation results for different scenarios is still an open problem. Existing approaches comprise the Pareto front or weighted sums. While the former is hardly informative for large-scale benchmarking (cf. §4), the latter depends on (arbitrarily) selected weights. In this work, we offer two alternative approaches while distinguishing two [^2] In reality, one can imagine a multitude of scenarios in between these two prototypical cases, hence we also consider benchmarking methods along this spectrum. What unites them, however, is their ability to aggregate multiple criteria. prototypical practical benchmarking scenarios with associated research questions (RQ):
Scenario 1 (Ranking).
First, consider a practitioner using open-ended text generation for a specific task, e.g., a customer support chatbot. This practitioner might primarily be interested in a complete scenario-specific relative ranking of existing methods. This motivation renders metric information about the methods’ performances a means to an end. Thus, an ordinal ranking of methods will do. RQ1: Can we exploit novel statistical methodologies for partial orders to establish multicriteria rankings that potentially allow for incomparability?
Scenario 2 (Cardinal Assessment).
Second, for researchers interested in designing new decoding methods (i.e., model, decoding strategy, or both), it is of utmost importance to know how much better one method is compared to another, i.e., having an absolute ranking on a cardinal scale. Knowledge of the performance of existing methods on different tasks will help derive new methods. RQ2: Can we aggregate multiple automatic evaluation metrics in a meaningful and statistically valid way?
Contributions.
We address RQ14) and RQ25) by proposing appropriate aggregation methods (cf. Fig. 1), including a novel summary metric to balance multiple assessments. We further provide experimental results by applying all introduced methods to over 1.8M stories generated by six LLMs on real-world datasets (cf. §3 for the setup and §4.2, §4.4, §5.2 for the results).
@@ 2 Related Work
Benchmarks are ubiquitous in applied machine learning (ML) research Zhang and Hardt (2024a); Shirali et al. (2023); Ott et al. (2022); Zhang et al. (2020); Thiyagalingam et al. (2022); Roelofs et al. (2019); Vanschoren et al. (2014), being used to make informed decisions and to demonstrate the superiority of newly proposed methods over concurrent ones Meyer et al. (2003); Hothorn et al. (2005); Eugster et al. (2012); Mersmann et al. (2015). In recent years, the focus has shifted towards multicriteria and multi-task benchmarking problems Cruz et al. (2024); Zhang and Hardt (2024b); Kohli et al. (2024); Jansen et al. (2024, 2023a, 2023b); Rodemann and Blocher (2024); Blocher et al. (2024). In a multitude of domains, there are several criteria concerning which methods need to be compared. Classical examples include runtime and accuracy in predictive ML Koch et al. (2015); Jansen et al. (2024) or performance and speed in optimization Schneider et al. (2018), to name only a few.
@@
Modern LLMs require evaluation across multiple metrics due to their broad capabilities (see, e.g., Wei et al., 2024; Liu et al., 2025). Assessing models on diverse tasks – from reasoning and comprehension to creativity and ethics – provides better understanding of their strengths and limitations Chiang et al. (2024). These comprehensive evaluation frameworks advance model performance while ensuring alignment with real-world applications and ethical standards Liu et al. (2023); Ji et al. (2023); Terry et al. (2023); Rodemann et al. (2025). Multicriteria benchmarking has thus become essential for guiding both theoretical progress and practical deployment of LLMs.
@@
Deepseek Wikitext Coherence Beam search $B$3, 5, 10, 15, 20, 50} 6 × 5261 × 6 = 189,396
@@
Llama3 Temperature sampling $\tau$0.1, 0.3, 0.5, 0.7, 0.9, 1.0} 6 × 5261 × 6 = 189,396
Mistralv03 Top-$k$ sampling $k$1, 3, 5, 10, 15, 20, 50} 6 x 5261 x 7 = 220,962
Qwen2 Top-$p$ (nucleus) sampling $p$0.6, 0.7, 0.8, 0.9, 0.95} 6 × 5261 × 5 = 157,830
@@
Table 1: Experimental setup: Over 1.8M text generations produced using various models and decoding strategies with different hyperparameter configurations. Prompts were drawn from three datasets (Wikitext, Wikinews, and Book), and outputs were evaluated on Coherence, Diversity, and Generation Perplexity.
Decoding methods for open-ended text generation are no exception. Several metrics to evaluate the quality of decoding strategies have been proposed and discussed in recent years Alihosseini et al. (2019); Celikyilmaz et al. (2021); Su and Xu (2022); Su et al. (2022); Gao et al. (2022); Becker et al. (2024); Garces-Arias et al. (2025). Diversity, MAUVE, coherence, and generation perplexity are among the most popular metrics. Diversity measures lexical variation using $n$-gram repetition rates, with higher scores indicating less repetition. MAUVE is a distribution similarity metric between generations and reference texts. Coherence is defined as the averaged log-likelihood of the generated text conditioned on the prompt and rewards logical flow. Finally, generation perplexity Jelinek et al. (2005) measures the predictability of the generated text under the language model; lower perplexity indicates that the text is more likely according to the model’s own probability distribution.
This multitude of quality metrics naturally raises the question of how to aggregate them, i.e., how to account for multiple dimensions of text quality to compare decoding methods holistically. It is self-evident that focusing on single metrics has obvious shortcomings. Exclusively optimizing for coherence will favor decoding methods with only moderate diversity, leading to degenerate, i.e., repetitive and uncreative generations (Holtzman et al., 2019; Lee et al., 2022). On the other hand, focusing solely on diversity will eventually result in incoherent text only slightly – if at all – related to the prompt. In this work, we offer a fresh perspective on the problem of multicriteria evaluation, adopting recent developments in the theory of depth functions and order theory (cf. §4).
3 Experimental Setup
@@
We evaluate six model architectures that generated over 1.8 million stories based on prompts sourced from three distinct datasets, utilizing five decoding strategies across 59 hyperparameter configurations.
Models.
@@
We employ GPT2-XL (1.5B, Radford et al., 2019), Mistral 7B v0.3 (Jiang et al., 2023, 2024), Llama 3.1 8B (Dubey et al., 2024), Deepseek 7B (DeepSeek-AI et al., 2024), Qwen 2 7B (Yang et al., 2024), and Falcon 2 11B (Malartic et al., 2024).
Evaluation Metrics.
Building upon Su and Collier (2023), we select diversity, coherence, and generation perplexity [^3] For their definitions, please refer to Appendix A. as automatic metrics to assess the quality of the generated texts individually. Based on this subset of possible instance-level metrics, we construct partial orders for multicriteria rankings (§4) and develop a cardinal assessment that collapses all metrics into one single score (§5). Since both approaches require instance-level metrics, we exclude MAUVE in this study as it assesses distributional similarities between samples of machine-generated text and human-written continuations, i.e. it relies on aggregated data, which would prevent us from applying the methods proposed in §4 and §5.
Datasets.
@@
We evaluate our methods across three domains for open-ended text generation: News, Wikipedia articles, and stories. Specifically, we use 2,000 articles from Wikinews for the news domain; 1,314 articles from the WikiText-103 dataset (Merity et al., 2016) for the Wikipedia domain; and 1,947 examples from the Project Gutenberg split of the BookCorpus (Zhu et al., 2015) for the story domain. Each example consists of a prompt and a gold reference (i.e., a human continuation) for evaluation. Further, we utilize the dataset provided by Garces-Arias et al. (2025), including over 1.8M generated continuations (with a maximum length of 256 tokens) for each prompt, along with aggregated metrics (coherence, diversity, MAUVE). We extend this dataset by computing sentence-level metrics and incorporating generation perplexity.
Decoding Strategies and Hyperparameters.
@@
For contrastive search (CS, Su et al., 2022), we evaluate 35 combinations of $\alpha$ and $k$, while for beam search (BS, Freitag and Al-Onaizan, 2017), we consider six beam widths $B$. For temperature sampling (Ackley et al., 1985), we consider six different temperatures $\tau$, for top-$k$ sampling (Fan et al., 2018), we use 7 different $k$ values and for top-$p$ (nucleus) sampling (Holtzman et al., 2019) we evaluate five different values for $p$, for a total of 59 decoding strategies configurations. All details are summarized in Table 1.
4 Scenario 1: Ranking Methods
To benchmark decoding methods according to multiple criteria (cf. §2) aiming for a ranking of methods (Scenario 1 and RQ1 in §1), we adopt very recent developments in the theory of multicriteria and multitask benchmarking Jansen et al. (2023b, a); Cruz et al. (2024); Zhang and Hardt (2024b); Kohli et al. (2024); Jansen et al. (2024); Rodemann and Blocher (2024); Blocher et al. (2024), some of them grounded in decision theory (social choice theory), some in the theory of data depth.
In this section, we propose benchmarking of decoding methods in terms of an ordinal ranking along (i) the extended Bradley-Terry model 4.1; Bradley and Terry, 1952b) and (ii) the union-free-generic (ufg) depth 4.3; Blocher et al., 2024; Blocher and Schollmeyer, 2024) as an alternative approach. Both approaches deliver ordinal rankings of decoding methods rather than a cardinal quality assessment (cf. left and middle column of Table 2). This can be motivated from a practical perspective (cf. §1): The cardinal information incorporated in numerous metrics can be considered redundant in cases when pure ranking of the decoding methods is the overall aim of benchmarking, not assigning scores to them. After all, a decoding method can either be deployed by practitioners or not, rendering the metric information not of primary practical interest.
Use Case
To illustrate our evaluation methodology, we apply it to the WikiText-103 dataset, which comprises 1,314 human-written prompts. We assess decoding methods by analyzing their text generations across three quality metrics: coherence, generation perplexity, and diversity. Our benchmarking approach produces partial rankings by determining whether one decoding method outperforms another, without quantifying the magnitude of performance differences.
[…]